OpenStudy (anonymous):

How to solve this inequalities problem? 1. -2x^2+6x+15≤0 2. 2x^3+13x^2-8x-46≥6 3. 5/x-6>3/x+2 4. 3x/x-1≤(x/x+4) +3

OpenStudy (anonymous):

which one you want first?

OpenStudy (anonymous):

lets start with this one\[-2x^2+6x+15≤0\]

OpenStudy (anonymous):

you have to factor it, i hate factoring, but it is easiest if the leading coefficient is positive, so lets change the signs and the inequality to get \[2x^2-6x-15\geq 0\] the factor as damn it doesn't factor you sure this is the exact problem? we can still solve it, it will just be more of a pain

OpenStudy (anonymous):

Yes that's the problem. Idk factoring im not good at it haha lol

OpenStudy (anonymous):

I know the rules but im still confuse with inequalities lol

OpenStudy (anonymous):

go to mathway.com

OpenStudy (anonymous):

you will get the exact answer if you type the inequalities in

OpenStudy (anonymous):

I don't have credit card. It doesn't show the solution only the answer

ganeshie8 (ganeshie8):

use quadratic formula and find x-intercepts

OpenStudy (anonymous):

How? :/

ganeshie8 (ganeshie8):

\(2x^2-6x-15 = 0 \)

ganeshie8 (ganeshie8):

solve x using quadratic formula

ganeshie8 (ganeshie8):

here is the formula :- \(x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):